REU 2008: SCHEDULE FOR WEEK EIGHT

Date  Babai May Postdocs* Student talks
Mon, August 11 TBA TBA  TBA  TBA
Tue, August 12 TBA TBA  TBA  TBA
Wed, August 13 TBA TBA  TBA  TBA
Thu, August 14 TBA TBA  TBA  TBA
Fri, August 15 TBA TBA  TBA  TBA

*Angeltveik, Basak, Crans, Schmidt, McReynolds

This year we have no work (YSP or SESAME) scheduled for the last week. There is time for many student presentations, and all are encouraged to speak. In addition, Professors Babai and May, and probably some of the postdocs and perhaps others will give special presentations, not necessarily connected with earlier talks.

And of course there will be food and parties.

Participants:

EIGHTH WEEK:

Babai: TBA May: TBA Postdocs: TBA